OpenStudy (anonymous):

What is the derivative of 2sqrtx?

OpenStudy (anonymous):

rewrite \(2\sqrt{x}\) using exponents then you can use the power rule.

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

Hi Pilot, 2x^1/2 Is the derivative just x?

OpenStudy (anonymous):

no, use the power rule...

OpenStudy (anonymous):

\[2x^{\frac{1}{2}}\]

OpenStudy (anonymous):

\[y=2x^{\frac{ 1 }{ 2 }} \Rightarrow \frac{ dy }{ dx }=2\cdot \frac{ 1 }{ 2 }\cdot x^{\left( \frac{ 1 }{ 2 }-1 \right)}\]

OpenStudy (anonymous):

can I multiply the 2 and 1/2 to get \[x^\frac{ -1 }{ 2 }\] as the derivative?